OpenStudy (hba):

Checking my answers if f(x)=x^3-9x^2+15x+3 What is the relative maxima ?

OpenStudy (hba):

I got 10.

OpenStudy (anonymous):

Really??

OpenStudy (hba):

Yeah really.

OpenStudy (anonymous):

It will be right as it is calculated by the Great @hba ha ha ha..

OpenStudy (hba):

Thanks.

OpenStudy (hba):

I have more question to get checked can i ask here or in a new post.

OpenStudy (anonymous):

I don't know how to solve these types of questions.. Can you teach me how to solve it..??

OpenStudy (hba):

@waterineyes Ha Ha Ha Nice joke,first let me ask my questions.

OpenStudy (anonymous):

First, we will take derivative.. Right??

OpenStudy (hba):

Yeah.

OpenStudy (anonymous):

So wait, I am doing it..

OpenStudy (hba):

okay :(

OpenStudy (anonymous):

f'(x) = 3x^2 - 18x + 15

OpenStudy (sirm3d):

there are two extrema, 1 relative maximum and 1 relative minimum.

OpenStudy (anonymous):

Should I factorize it??

OpenStudy (hba):

put f'(x)=0

OpenStudy (anonymous):

Yes I mean the same..

OpenStudy (hba):

Then the answer is yes

OpenStudy (anonymous):

\[x^2 - 6x + 5 = 0 \implies (x - 5)(x - 1) = 0\]

OpenStudy (anonymous):

x = 5 and x = 1

OpenStudy (hba):

Yes you are right

OpenStudy (hba):

Now put 1 in the orignal equation.

OpenStudy (anonymous):

Why not 5 ??

OpenStudy (hba):

Because we are finding the relative maxima

OpenStudy (anonymous):

f(1) = 10

OpenStudy (sirm3d):

oooh, i misread the problem. relative maxima.

OpenStudy (hba):

Correct.

OpenStudy (anonymous):

So, we have to put smaller value??

OpenStudy (hba):

Yes.

OpenStudy (anonymous):

Just checking your answer.. Ha ha ha..

OpenStudy (anonymous):

Now don't say, to find relative minima, we will put 5..

OpenStudy (hba):



OpenStudy (hba):

@waterineyes Check these two questions.

OpenStudy (anonymous):

How to solve these now??

OpenStudy (hba):

I got f'(x)=e in both the cases.

OpenStudy (hba):

And f''(x)=1/e for 1st and f''(x)=e in the 2nd one.

OpenStudy (hba):

So i am right ?

OpenStudy (anonymous):

Are you sure it is e ??

OpenStudy (hba):

Pretty sure Because i did this one like 100 times when i was in High school

OpenStudy (anonymous):

You have used division rule here??

OpenStudy (hba):

U/V form.

OpenStudy (anonymous):

For first : \[f'(x) = \frac{1 - \ln(x)}{x^2}\]

OpenStudy (hba):

Wait lemme check.

OpenStudy (anonymous):

Take your time..

OpenStudy (hba):

Yeah right

OpenStudy (hba):

@waterineyes ?

OpenStudy (anonymous):

Who is right??

OpenStudy (hba):

The most intelligent @waterineyes

OpenStudy (anonymous):

Not intelligent as you @hba

OpenStudy (anonymous):

Now, find the second derivative..

OpenStudy (hba):

Ha Ha Ha Yeah wait.

OpenStudy (hba):

I am getting something,which looks extremly wrong :/